# The complex structure of the Teichmüller space of circle diffeomorphisms in the Zygmund smooth class

### v4

@@ 2. The complex Banach structure
The existence of a local holomorphic right inverse $s_{\nu}$ of $\Phi$ such that $s_{\nu}\circ\Phi(\nu)=\nu$ can be proved by the standard argument if $f^{\nu}$ is bi-Lipschitz in the hyperbolic metric for $\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$. This is obtained as the generalized Ahlfors–Weill section by using the bi-Lipschitz quasiconformal reflection with respect to the quasicircle $f_{\nu}(\mathbb{S})$. See [5, 11] for a general argument, and [10, 16] for its application to particular Teichmüller spaces. In our present case, the same proof can be applied.
For an arbitrary $\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$, Lemma 2.3 shows that there exists $\nu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ such that $\nu$ is Teichmüller equivalent to $\mu$ and $f^{\nu}$ is bi-Lipschitz in the hyperbolic metric. Then, we can take a local holomorphic right inverse $s_{\nu}$ defined on some neighborhood $V_{\varphi}$ of $\varphi=\Phi(\nu)$ in $A^{Z}(\mathbb{D})$ such that $s_{\nu}\circ\Phi(\mu)=s_{\nu}\circ\Phi(\nu)=\nu$. By using the right translations $r_{\mu}$ and $r_{\nu}$ which are biholomorphic automorphisms of ${\rm Bel}^{Z}(\mathbb{D}^{*})$ as in Lemma 2.2, we set $s_{\mu}=r_{\mu}^{-1}\circ r_{\nu}\circ s_{\nu}$. This is a local holomorphic right inverse of $\Phi$ defined on $V_{\varphi}$ satisfying $s_{\mu}\circ\Phi(\mu)=\mu$. ∎
The statements of Theorem 1.2 correspond to the following corollaries to Theorem 2.4.
Corollary 2.5.
The Schwarzian derivative map $\Phi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to A^{Z}(\mathbb{D})$ is a holomorphic split submersion onto the image.
@@
By Theorem 1.1, $\Phi$ is holomorphic, and by Theorem 2.4, it is a holomorphic split submersion. ∎
@@
For a surjective holomorphic map $\Phi:B\to A$ from a domain $B\subset Y$ to a domain $A\subset X$ of Banach spaces $X$ and $Y$ in general, $\Phi$ is a holomorphic split submersion if and only if both of the following conditions are satisfied (see [12, p.89]):
@@
For certain Teichmüller spaces defined by the supremum norm such as the universal Teichmüller space and Teichmüller spaces of circle diffeomorphisms, it has been proved that the Schwarzian derivative map $\Phi$ is a holomorphic split submersion. However, for Teichmüller spaces defined by the integrable norm such as the Weil–Petersson Teichmüller space and the BMO Teichmüller space, only condition (1) has been verified.
Corollary 2.6.
The Bers embedding $\alpha:T^{Z}\to A^{Z}(\mathbb{D})$ is a homeomorphism onto the image $\alpha(T^{Z})=\Phi({\rm Bel}^{Z}(\mathbb{D}^{*}))$, which is a connected open subset of $A^{Z}(\mathbb{D})$.
@@
The image $\alpha(T^{Z})=\Phi({\rm Bel}^{Z}(\mathbb{D}^{*}))$ is open because for every $\varphi\in\Phi({\rm Bel}^{Z}(\mathbb{D}^{*}))$, the neighborhood $V_{\varphi}$ as in Theorem 2.4 is contained in $\Phi({\rm Bel}^{Z}(\mathbb{D}^{*}))$. By Theorem 1.1, the Bers embedding $\alpha:T^{Z}\to\alpha(T^{Z})$ is continuous. Conversely, since there is a local holomorphic right inverse $s_{\mu}$ of $\Phi$ defined on some neighborhood $V_{\varphi}$ for every $\varphi\in\alpha(T^{Z})$ satisfying $\alpha^{-1}|_{V_{\varphi}}=\pi\circ s_{\mu}$ with $\Phi(\mu)=\varphi$, we see that $\alpha^{-1}$ is continuous. ∎
Corollary 2.7.
The Teichmüller space $T^{Z}$ is endowed with the unique complex Banach manifold structure such that the Teichmüller projection $\pi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to T^{Z}$ is a holomorphic map with surjective derivatives at all points of ${\rm Bel}^{Z}(\mathbb{D}^{*})$.
@@
As the Bers embedding $\alpha:T^{Z}\to A^{Z}(\mathbb{D})$ is a homeomorphism onto the image by Corollary 2.6, $T^{Z}$ is endowed with the complex Banach structure as the domain $\alpha(T^{Z})$ in $A^{Z}(\mathbb{D})$. The uniqueness follows from [4, Proposition 8]. ∎
@@ 3. Pre-Schwarzian derivatives
Let $\psi=\log f^{\prime}$ for a locally univalent holomorphic function $f$ on $\mathbb{D}$. Its derivative $P_{f}=\psi^{\prime}$ is called the pre-Schwarzian derivative of $f$. Let $B(\mathbb{D})$ be the space of all holomorphic functions $\psi$ on $\mathbb{D}$ such that
$\|\psi\|_{B}=\sup_{|z|<1}\,(1-|z|^{2})|\psi^{\prime}(z)|<\infty.$
@@
Such a $\psi$ is called a Bloch function. Then, ignoring the difference of complex constant functions, we can regard $B(\mathbb{D})$ as the complex Banach space with norm $\|\cdot\|_{B}$. We note that $\psi\in B(\mathbb{D})$ if and only if $\psi^{\prime}\in A^{Z}(\mathbb{D})$.
@@
We require that $f_{\mu}$ satisfies $f_{\mu}(\infty)=\infty$. Then, $P_{{f_{\mu}}|_{\mathbb{D}}}$ is well defined for every $\mu\in{\rm Bel}(\mathbb{D}^{*})$ by this normalization and $\psi=\log\,({f_{\mu}}|_{\mathbb{D}})^{\prime}$ belongs to $B(\mathbb{D})$. We may assume that $\psi(0)=0$. By this correspondence $\mu\mapsto\log\,({f_{\mu}}|_{\mathbb{D}})^{\prime}$, we have a holomorphic map $\Psi:{\rm Bel}(\mathbb{D}^{*})\to B(\mathbb{D})$ called the pre-Schwarzian derivative map.
@@
Let $\widetilde{\mathcal{T}}=\Psi({\rm Bel}(\mathbb{D}^{*}))$, which consists of those $\psi=\log f^{\prime}$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with $f(\mathbb{D})$ bounded. It is known that this is an open subset of $B(\mathbb{D})$ but there are other uncountably many connected components of the open subset in $B(\mathbb{D})$ consisting of those $\psi=\log f^{\prime}$ with $f(\mathbb{D})$ unbounded. See [17].
In the case of the Teichmüller space $T^{Z}$, we define the corresponding space
$B^{Z}(\mathbb{D})=\{\psi\in B(\mathbb{D})\mid\|\psi\|_{B^{Z}}=|\psi^{\prime}(0)|+\sup_{|z|<1}\,(1-|z|^{2})|\psi^{\prime\prime}(z)|<\infty\},$
@@
which we regard as the complex Banach space by ignoring the difference of complex constant functions. We note that $\psi\in B^{Z}(\mathbb{D})$ if and only if $\psi^{\prime}\in B(\mathbb{D})$. Moreover, we see that $\|\psi\|_{B}\lesssim\|\psi\|_{B^{Z}}$ (see [19, Theorem 5.4]). Hereafter, the symbol $\lesssim$ stands for the inequality modulo a uniform positive constant multiple. The space $B^{Z}(\mathbb{D})$ was defined in [14] by using $\sup_{z\neq w\in\mathbb{D}}|\psi^{\prime}(z)-\psi^{\prime}(w)|/d(z,w)$ for the hyperbolic distance $d$ on $\mathbb{D}$, but they are equivalent (see [19, Theorem 5.5]).
@@
Concerning the pre-Schwarzian derivative model of $T^{Z}$, the following theorem has been obtained:
Theorem 3.1 ([14, Theorem 1.4]).
The image of ${\rm Bel}^{Z}(\mathbb{D}^{*})$ under the pre-Schwarzian derivative map $\Psi$ is contained in $B^{Z}(\mathbb{D})$ and $\Psi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to B^{Z}(\mathbb{D})$ is holomorphic.
@@
We set this image as $\widetilde{\mathcal{T}}^{Z}=\Psi({\rm Bel}^{Z}(\mathbb{D}^{*}))\subset B^{Z}(\mathbb{D})$. It can be seen that
$\widetilde{\mathcal{T}}^{Z}=B^{Z}(\mathbb{D})\cap\widetilde{\mathcal{T}}$
@@
by [14, Theorem 1.1]. This is an open subset of $B^{Z}(\mathbb{D})$ because $\widetilde{\mathcal{T}}$ is open in $B(\mathbb{D})$ and $\|\!\!\!~{}\cdot~{}\!\!\|_{B}\lesssim\|\cdot\|_{B^{Z}}$. Moreover, unlike in the case of the universal Teichmüller space $T$, it is proved in [14, Theorem 1.3] that $\widetilde{\mathcal{T}}^{Z}$ is exactly the set of those $\psi=\log f^{\prime}$ in $B^{Z}(\mathbb{D})$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with its complex dilatation in ${\rm Bel}^{Z}(\mathbb{D}^{*})$. In other words, for every $f=f_{\mu}|_{\mathbb{D}}$ with $\mu\in{\rm Bel}^{Z}(\mathbb{D}^{*})$ and $\log f^{\prime}\in B^{Z}(\mathbb{D})$, its image $f(\mathbb{D})$ is a bounded domain in $\mathbb{C}$.
@@
We examine the structure of $\widetilde{\mathcal{T}}^{Z}$ and the pre-Schwarzian derivative map $\Psi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to\widetilde{\mathcal{T}}^{Z}$. The strategy is to factorize the Schwarzian derivative map $\Phi$ by $\Psi$ and to bring the properties of $\Phi$ to $\Psi$. Due to the relation $S_{f}=(P_{f})^{\prime}-(P_{f})^{2}/2$, we consider the following map $\Lambda$ satisfying $\Lambda\circ\Psi=\Phi$.
Lemma 3.2.
The map $\Lambda:B^{Z}(\mathbb{D})\to A^{Z}(\mathbb{D})$ defined by $\Lambda(\psi)=\psi^{\prime\prime}-(\psi^{\prime})^{2}/2$ is holomorphic.
@@
It is easy to see that $\Lambda$ is Gâteaux holomorphic. Hence, to prove that $\Lambda$ is holomorphic, it suffices to show that $\Lambda$ is locally bounded. See [1, p.28] and [2, Theorem 14.9]. The local boundedness can be verified as follows.
Let $\psi\in B^{Z}(\mathbb{D})$. Then $\psi^{\prime\prime}\in A^{Z}(\mathbb{D})$ with $\|\psi^{\prime\prime}\|_{A^{Z}}\leq\|\psi\|_{B^{Z}}$ by definition. Moreover, we have $\|(\psi^{\prime})^{2}\|_{A^{Z}}\lesssim\|\psi\|_{B^{Z}}^{2}$. Indeed,
$\|(\psi^{\prime})^{2}\|_{A^{Z}}\leq\left(\sup_{|z|<1}\,(1-|z|^{2})^{1/2}|\psi^{\prime}(z)|\right)^{2},$
and, by [18, Proposition 8], we see that
$\sup_{|z|<1}\,(1-|z|^{2})^{1/2}|\psi^{\prime}(z)|\asymp|\psi^{\prime}(0)|+\sup_{|z|<1}\,(1-|z|^{2})^{3/2}|\psi^{\prime\prime}(z)|\leq\|\psi\|_{B^{Z}}.$
Therefore,
$\|\Lambda(\psi)\|_{A^{Z}}\leq\|\psi^{\prime\prime}\|_{A^{Z}}+\frac{1}{2}\|(\psi^{\prime})^{2}\|_{A^{Z}}\lesssim\|\psi\|_{B^{Z}}+\frac{1}{2}\|\psi\|_{B^{Z}}^{2},$
@@
We can improve Theorem 3.1 immediately in virtue of Theorem 2.4. It also asserts the same property for $\Lambda=\Lambda|_{\widetilde{\mathcal{T}}^{Z}}$.
Theorem 3.3.
Both $\Psi:{\rm Bel}^{Z}(\mathbb{D}^{*})\to\widetilde{\mathcal{T}}^{Z}\subset B^{Z}(\mathbb{D})$ and $\Lambda:\widetilde{\mathcal{T}}^{Z}\to\alpha(T^{Z})\subset A^{Z}(\mathbb{D})$ are holomorphic split submersions.
@@
By Theorem 2.4, $\Phi=\Lambda\circ\Psi$ is a holomorphic split submersion. Combined with Theorem 3.1 and Lemma 3.2, this implies the statement. ∎

### v5

@@ 2. The complex Banach structure
The existence of a local holomorphic right inverse $s_{\nu}$ of $\Phi$ such that $s_{\nu}\circ\Phi(\nu)=\nu$ can be proved by the standard argument if $f^{\nu}$ is bi-Lipschitz in the hyperbolic metric for $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$. This is obtained as the generalized Ahlfors–Weill section by using the bi-Lipschitz quasiconformal reflection with respect to the quasicircle $f_{\nu}(\mathbb{S})$. See [3] for a general argument, and [6, 12] for its application to particular Teichmüller spaces. In our present case, the same proof can be applied.
For an arbitrary $\mu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$, Lemma 6 shows that there exists $\nu\in{\rm Bel}_{Z}(\mathbb{D}^{*})$ such that $\nu$ is Teichmüller equivalent to $\mu$ and $f^{\nu}$ is bi-Lipschitz in the hyperbolic metric. Then, we can take a local holomorphic right inverse $s_{\nu}$ defined on some neighborhood $V_{\varphi}$ of $\varphi=\Phi(\nu)$ in $A_{Z}(\mathbb{D})$ such that $s_{\nu}\circ\Phi(\mu)=s_{\nu}\circ\Phi(\nu)=\nu$. By using the right translations $r_{\mu}$ and $r_{\nu}$ which are biholomorphic automorphisms of ${\rm Bel}_{Z}(\mathbb{D}^{*})$ as in Lemma 5, we set $s_{\mu}=r_{\mu}^{-1}\circ r_{\nu}\circ s_{\nu}$. This is a local holomorphic right inverse of $\Phi$ defined on $V_{\varphi}$ satisfying $s_{\mu}\circ\Phi(\mu)=\mu$. ∎
The statements of Theorem 3 correspond to the following corollaries to Theorem 7.
Corollary 8.
The Schwarzian derivative map $\Phi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to A_{Z}(\mathbb{D})$ is a holomorphic split submersion.
@@
By Theorem 1, $\Phi$ is holomorphic, and by Theorem 7, it is a holomorphic split submersion. ∎
@@
For a surjective holomorphic map $\Phi:B\to A$ from a domain $B\subset Y$ to a domain $A\subset X$ of Banach spaces $X$ and $Y$ in general, $\Phi$ is a holomorphic split submersion if and only if both of the following conditions are satisfied (see [8, p.89]):
@@
For certain Teichmüller spaces defined by the supremum norm such as the universal Teichmüller space and Teichmüller spaces of circle diffeomorphisms, it has been proved that the Schwarzian derivative map $\Phi$ is a holomorphic split submersion. However, for Teichmüller spaces defined by the integrable norm, only condition (1) has been verified.
Corollary 9.
The Bers embedding $\alpha:T_{Z}\to A_{Z}(\mathbb{D})$ is a homeomorphism onto the image $\alpha(T_{Z})=\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$, which is a connected open subset of $A_{Z}(\mathbb{D})$.
@@
The image $\alpha(T_{Z})=\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$ is open because for every $\varphi\in\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$, the neighborhood $V_{\varphi}$ as in Theorem 7 is contained in $\Phi({\rm Bel}_{Z}(\mathbb{D}^{*}))$. By Corollary 2, the Bers embedding $\alpha:T_{Z}\to\alpha(T_{Z})$ is continuous. Conversely, since there is a local holomorphic right inverse $s_{\mu}$ of $\Phi$ defined on some neighborhood $V_{\varphi}$ for every $\varphi\in\alpha(T_{Z})$ satisfying $\alpha^{-1}|_{V_{\varphi}}=\pi\circ s_{\mu}$ with $\Phi(\mu)=\varphi$, we see that $\alpha^{-1}$ is continuous. ∎
Corollary 10.
The Teichmüller space $T_{Z}$ is endowed with the unique complex Banach manifold structure such that the Teichmüller projection $\pi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to T_{Z}$ is a holomorphic map with surjective derivatives at all points of ${\rm Bel}_{Z}(\mathbb{D}^{*})$.
@@
As the Bers embedding $\alpha:T_{Z}\to A_{Z}(\mathbb{D})$ is a homeomorphism onto the image by Corollary 9, $T_{Z}$ is endowed with the complex Banach structure as the domain $\alpha(T_{Z})$ in $A_{Z}(\mathbb{D})$. The uniqueness follows from [2, Proposition 8]. ∎
@@ 3. Pre-Schwarzian derivatives
Let $\psi_{f}(z)=\log f^{\prime}(z)$ for a locally univalent holomorphic function $f$ on $\mathbb{D}$. Its derivative $P_{f}(z)=\psi_{f}^{\prime}(z)$ is called the pre-Schwarzian derivative of $f$. Let $B(\mathbb{D})$ be the space of all holomorphic functions $\psi$ on $\mathbb{D}$ such that
$\|\psi\|_{B}={\rm sup}_{|z|<1}(1-|z|^{2})|\psi^{\prime}(z)|<\infty.$
@@
Such a $\psi$ is called a Bloch function. Then, ignoring the difference of complex constant functions, we can regard $B(\mathbb{D})$ as the complex Banach space with norm $\|\cdot\|_{B}$. We note that $\psi\in B(\mathbb{D})$ if and only if $\psi^{\prime}\in A_{Z}(\mathbb{D})$.
@@
We require that $f_{\mu}$ satisfies $f_{\mu}(\infty)=\infty$. Then, $P_{{f_{\mu}}|_{\mathbb{D}}}$ is well defined for every $\mu\in{\rm Bel}(\mathbb{D}^{*})$ by this normalization and $\psi_{{f_{\mu}}|_{\mathbb{D}}}$ belongs to $B(\mathbb{D})$. We may assume that $\psi_{{f_{\mu}}|_{\mathbb{D}}}(0)=0$. By this correspondence $\mu\mapsto\psi_{{f_{\mu}}|_{\mathbb{D}}}$, we have a holomorphic map $\Psi:{\rm Bel}(\mathbb{D}^{*})\to B(\mathbb{D})$ called the pre-Schwarzian derivative map.
@@
Let $\widetilde{\mathcal{T}}=\Psi({\rm Bel}(\mathbb{D}^{*}))$, which consists of those $\psi_{f}$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with $f(\mathbb{D})$ bounded. It is known that this is an open subset of $B(\mathbb{D})$ but there are other uncountably many connected components of the open subset in $B(\mathbb{D})$ consisting of those $\psi_{f}$ with $f(\mathbb{D})$ unbounded. See [13].
In the case of the Teichmüller space $T_{Z}$, we define the corresponding space
$B_{Z}(\mathbb{D})=\{\psi\in B(\mathbb{D})\mid\|\psi\|_{B_{Z}}=|\psi^{\prime}(0)|+{\rm sup}_{|z|<1}(1-|z|^{2})|\psi^{\prime\prime}(z)|<\infty\},$
@@
which we regard as the complex Banach space by ignoring the difference of complex constant functions. We note that $\psi\in B_{Z}(\mathbb{D})$ if and only if $\psi^{\prime}\in B(\mathbb{D})$. Moreover, we see that $\|\psi\|_{B}\lesssim\|\psi\|_{B_{Z}}$ (see [15, Theorem 5.4]). Hereafter, the symbol $\lesssim$ stands for the inequality modulo a uniform positive constant multiple. The space $B_{Z}$ was defined in [10] by using $\sup_{z\neq w\in\mathbb{D}}|\psi^{\prime}(z)-\psi^{\prime}(w)|/d(z,w)$ for the hyperbolic distance $d$ on $\mathbb{D}$, but they are equivalent (see [15, Theorem 5.5]).
@@
Concerning the pre-Schwarzian derivative model of $T_{Z}$, the following theorem has been obtained:
Theorem 11 ([10, Theorem 1.4]).
The image of ${\rm Bel}_{Z}(\mathbb{D}^{*})$ under the pre-Schwarzian derivative map $\Psi$ is contained in $B_{Z}(\mathbb{D})$ and $\Psi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to B_{Z}(\mathbb{D})$ is holomorphic.
@@
We set this image as $\widetilde{\mathcal{T}}_{Z}=\Psi({\rm Bel}_{Z}(\mathbb{D}^{*}))\subset B_{Z}(\mathbb{D})$. It can be seen that
$\widetilde{\mathcal{T}}_{Z}=B_{Z}(\mathbb{D})\cap\widetilde{\mathcal{T}}$
@@
by [10, Theorem 1.1]. This is an open subset of $B_{Z}(\mathbb{D})$ because $\widetilde{\mathcal{T}}$ is open in $B(\mathbb{D})$ and $\|\cdot\|_{B}\lesssim\|\cdot\|_{B_{Z}}$. Moreover, differently from the case of the universal Teichmüller space $T$, it is proved in [10, Theorem 1.3] that $\widetilde{\mathcal{T}}_{Z}$ is exactly the set of those $\psi_{f}$ in $B_{Z}(\mathbb{D})$ for a conformal homeomorphism $f$ of $\mathbb{D}$ that is quasiconformally extendable to $\mathbb{D}^{*}$ with its complex dilatation in ${\rm Bel}_{Z}(\mathbb{D}^{*})$.
@@
We examine the structure of $\widetilde{\mathcal{T}}_{Z}$ and the pre-Schwarzian derivative map $\Psi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to\widetilde{\mathcal{T}}_{Z}$. The strategy is to factorize the Schwarzian derivative map $\Phi$ by $\Psi$ and to bring the properties of $\Phi$ to $\Psi$. Due to the relation $S_{f}=(P_{f})^{\prime}-(P_{f})^{2}/2$, we consider the following map $\Lambda$ satisfying $\Lambda\circ\Psi=\Phi$.
Lemma 12.
The map $\Lambda:B_{Z}(\mathbb{D})\to A_{Z}(\mathbb{D})$ defined by $\Lambda(\psi)=\psi^{\prime\prime}-\frac{1}{2}(\psi^{\prime})^{2}$ is holomorphic.
@@
It is easy to see that $\Lambda$ is Gâteaux holomorphic. Hence, to prove that $\Lambda$ is holomorphic, it suffices to show that $\Lambda$ is locally bounded. See [1, p.28]. The local boundedness can be verified as follows.
Let $\psi\in B_{Z}(\mathbb{D})$. Then $\psi^{\prime\prime}\in A_{Z}(\mathbb{D})$ with $\|\psi^{\prime\prime}\|_{A_{Z}}\leq\|\psi\|_{B_{Z}}$ by definition. Moreover, for any $\phi\in B_{Z}(\mathbb{D})$, we have $\phi^{\prime}\psi^{\prime}\in A_{Z}(\mathbb{D})$ with $\|\phi^{\prime}\psi^{\prime}\|_{A_{Z}}\lesssim\|\phi\|_{B_{Z}}\|\psi\|_{B_{Z}}$, and in particular, $\|(\psi^{\prime})^{2}\|_{A_{Z}}\lesssim\|\psi\|_{B_{Z}}^{2}$. Indeed,
$\|\phi^{\prime}\psi^{\prime}\|_{A_{Z}}=\sup_{|z|<1}(1-|z|^{2})|\phi^{\prime}(z)||\psi^{\prime}(z)|\leq\sup_{|z|<1}(1-|z|^{2})^{1/2}|\phi^{\prime}(z)|\cdot\sup_{|z|<1}(1-|z|^{2})^{1/2}|\psi^{\prime}(z)|.$
Here, by [14, Proposition 8], we see that
$\sup_{|z|<1}(1-|z|^{2})^{1/2}|\psi^{\prime}(z)|\asymp|\psi^{\prime}(0)|+\sup_{|z|<1}(1-|z|^{2})^{3/2}|\psi^{\prime\prime}(z)|\leq\|\psi\|_{B_{Z}}.$
Thus, $\|\phi^{\prime}\psi^{\prime}\|_{A_{Z}}$ is bounded as required. Therefore,
$\|\Lambda(\psi)\|_{A_{Z}}\leq\|\psi^{\prime\prime}\|_{A_{Z}}+\frac{1}{2}\|(\psi^{\prime})^{2}\|_{A_{Z}}\lesssim\|\psi\|_{B_{Z}}+\frac{1}{2}\|\psi\|_{B_{Z}}^{2},$
@@
We state the main claim in this section, which improves Theorem 11.
Theorem 13.
Both $\Psi:{\rm Bel}_{Z}(\mathbb{D}^{*})\to\widetilde{\mathcal{T}}_{Z}\subset B_{Z}(\mathbb{D})$ and $\Lambda|_{\widetilde{\mathcal{T}}_{Z}}:\widetilde{\mathcal{T}}_{Z}\to\alpha(T_{Z})\subset A_{Z}(\mathbb{D})$ are holomorphic split submersions.
@@
By Theorem 7, $\Phi=\Lambda\circ\Psi$ is a holomorphic split submersion. Combined with Theorem 11 and Lemma 12, this implies the statement. ∎
In fact, $\widetilde{\mathcal{T}}_{Z}$ is a disk-bundle over $T_{Z}\cong\alpha(T_{Z})$.